Determine the resultant of two vectors using the component method and illustrate. Use separate sheet of paper and scientific calculator.


Given: A= 15 Units, 25 Counterclockwise (+) x-axis

B = 30 Units, 40 Counterclockwise (-) x-axis



1
Expert's answer
2021-11-07T19:26:04-0500
"A_x=15\cos25°=13.6,\\
A_y=15\sin25°=6.34.\\
B_x=-30\cos40°=-23,\\
B_y=-30\sin40°=-19.\\
R_x=A_x+B_x,\\
R_y=A_y+B_y,\\
R=\sqrt{R_x^2+R_y^2}=15.8,\\\space\\
\theta=\arctan\frac{R_y}{R_x}=53°\text{ counterclockwise from (-x)}"